Problem 2 (15 points) FIN300 Corporations 5-year bonds yield 7.00%, and 5-year T-bonds yield 5.15%. The real risk-free rate is r*-3.0%, the inflation premium for all 5-year bonds is IP 1.75%, the liquidity premium for FIN300s bonds is LP-0.75% versus zero for T-bonds, and the maturity risk premium for all bonds is found with the formula MRP-(-1) x 0.1%, where t number of years to maturity. What is the default risk premium (DRP) on FIN300s bonds?Can you please include the formula with the answer as well.
Answer #1

Solution-

5 year bond yield (r)= 7.0%

5 year T bond yield (rt) =5.15%

Real risk free rate (r*)= 3.0%

Liquidity Premium (LP) = 0.75%

Inflation premium (IP)= 1.75%

Market risk premium (MRP) =(t-1)*0.1%

MRP = (5-1)*0.1%= 0.40%

Default risk premium (DRP) =?

r= r*+DRP+LP+IP+MRP

7.0%= 3%+DRP+0.75%+1.75%+0.40%

DRP = 7.0%-3.0%-0.75%-1.75%-0.40%= 1.1%

Default risk premium (DRP ) = 1.1%